OpenStudy (anonymous):

2/3 (r - 2) - 1/2(r + 4) how does it equal 1/6r - 10/3

OpenStudy (anonymous):

=2/3r -4/3 -1/2r -2 =4/6r-3/6r -4/3 -6/3 =1/6r -10/3

OpenStudy (anonymous):

Just multiply, then sum

OpenStudy (anonymous):

thanks =)